Use the definition of the fact that f(x) is O(g(x)) to show that:

1. 7x^2 is O(x^2)

2. x^4 + 9x^3 + 4x + 7 is O(x^4)

3. (x^2+1) / (x+1) is O(x) Hint: Try to simplify algebraic expression.
Answer #1

`Hey,

Note: If you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

1)

Let f(x)=7*x^2

So,

f(x)<=8*x^2 for all x>=1

So, we found c=8 and x0=1

So,

f(x)=O(x^2)

2)

Let f(x)=x^4 + 9x^3 + 4x + 7

So,

f(x)=x^4 + 9x^3 + 4x + 7<=x^4+9*x^4+4*x^4+7*x^4=21*x^4 for x>=1

So, we found a constant c=21 and x0=1

So,

f(x)=O(x^4)

Note: Brother According to Chegg's policy we are only allowed to answer first 2 part if there are many. So, I request you to post other part as separate posts.

Kindly revert for any queries

Thanks.